HanAkoSolo (jamierox4ev3r):

Math Review Day 2! 6a. Rationalize the expression and simplify

HanAkoSolo (jamierox4ev3r):

Give me a second while I type out the expression. Thank you

HanAkoSolo (jamierox4ev3r):

\(\huge\frac{\sqrt{10}}{\sqrt{5}-2}\)

OpenStudy (anonymous):

do you know the root of ten?

HanAkoSolo (jamierox4ev3r):

yes. The root of 10 is not rational, so I don't see how knowing that would help me solve this problem

HanAkoSolo (jamierox4ev3r):

\(\huge\sqrt{10} = 3.16227766...\)

HanAkoSolo (jamierox4ev3r):

I have a very limited recollection on rationalizing. So an explanation on that would be nice as well, thank you

OpenStudy (anonymous):

i learned that you can use a perfect number to minimalize something like this heres an example the root of 18 equals 3root2

OpenStudy (nincompoop):

do you want to see if you can do anything by multiplying it by a conjugate

HanAkoSolo (jamierox4ev3r):

a conjugate...so like the opposite? do you mean multiplying by \(\huge\sqrt{5}+2\) ??

HanAkoSolo (jamierox4ev3r):

o-o

OpenStudy (anonymous):